Hydrogen gas at 1 atm pressure is bubbled through Ni2+ ions present in a purified solution
(obtained after solvent extraction). If the activity of Ni2+ in solution is 0.3, determine the minimum
pH required to precipitate Ni from the solution at 25oC. {Data: Standard oxidation potential for
H2/2H+ and Ni/Ni2+ are 0.0 V and 0.25 V respectively, R = 8.314 J/mol/K, Faraday constant = 96500
Coulomb/gm eqv wt}.
1
Expert's answer
2020-11-23T06:49:45-0500

E(Ni2+/Ni) = -0.25 + RT/(2F)*ln(a(Ni2+)) = -0.25+8.314*(25+273)/(2*96500)*ln(0.3) = -0.265

E(Ni2+/Ni) = E(H+/H2)

E(H+/H2) = RT/F*ln(a(H+)/p(H2)0.5)

-0.265 = 8.314*(25+273)/96500*(-2.30*pH)

pH = 4.49